# Benchmark Meta2 Evaluation - How to get it right

### v1

@@ Abstract
Recent advancements in Language Models (LMs) have catalyzed the creation of multiple benchmarks, designed to assess these models’ general capabilities.
A crucial task, however, is assessing the validity of the benchmarks themselves.
This is most commonly done via Benchmark Agreement Testing (BAT), where new benchmarks are validated against established ones using some agreement metric (e.g., rank correlation).
Despite the crucial role of BAT for benchmark builders and consumers,
there are no standardized procedures for such agreement testing.
This deficiency can lead to invalid conclusions, fostering mistrust in benchmarks and upending the ability to properly choose the appropriate benchmark to use.
By analyzing over 40 prominent benchmarks, we demonstrate how some overlooked methodological choices can significantly influence BAT results, potentially undermining the validity of conclusions.
To address these inconsistencies, we propose a set of best practices for BAT and demonstrate how utilizing these methodologies greatly improves BAT robustness and validity.
[…]
Our findings underscore the necessity for standardized BAT, ensuring the robustness and validity of benchmark evaluations in the evolving landscape of language model research.
[…]
Leaderboard: https://huggingface.co/spaces/per/BenchBench
@@ 1 Introduction
As Language Models (LMs) increasingly excel across a broad range of tasks, new benchmarks – often measuring similar abilities – are constantly proposed. This deluge in benchmarks highlights the role of Benchmark Agreement Testing (BAT); namely, validating some benchmarks using other benchmarks by measuring an agreement metric (e.g., Pearson correlation) over their models’ scores (Liu2021DoQA).
BAT is often used to validate that a new proposed benchmark measures what it was designed to measure. The expectations from this measurement depend on the benchmark’s goal; demonstrating high agreement can serve to show that a new benchmark captures similar model abilities to established benchmarks (Lei2023S3EvalAS; Viswanathan2023Prompt2ModelGD; Chang2023DoLM; Li2024TreeEvalBE; Prabhu2024LifelongBE; He2024UltraEvalAL). High agreement can also validate that an efficient version of a benchmark (e.g., requiring less compute or labeling) measures the same thing as the original benchmark (Perlitz2023EfficientB; Polo2024tinyBenchmarksEL; Prabhu2024LifelongBE; Vivek2023AnchorPB). In contrast, if a benchmark aims to test a unique trait – one that is not properly covered by existing benchmarks – BAT will be used to demonstrate the disagreement of such benchmarks with existing ones (Yuan2024PRobELMPR; waldis2024holmes). The above goals are relevant both for benchmark creators and for benchmark consumers. Creators will typically use BAT to validate the properties of their new benchmark; benchmark consumers might use it to choose which existing benchmark they want to use.
However, despite the wide application of BAT in recent years, there is a glaring absence of common methodology. Specifically, the significance of several BAT methodological decisions is currently overlooked, undermining the validity of any conclusions made.
@@
We focus on three such critical choices: selecting the reference benchmark (§3.1), the models included in the test (§3.2), as well as the correlation metrics and their interpretation (§3.3). For example, as seen in Figure 1, choosing a different subset of models produces substantially different correlation scores, leading to different conclusions about benchmark agreement. The figure exemplifies that two benchmarks can (and often do) show high agreement over a wide range of models, while agreement over a few top-ranked models remains low.
[…]
@@
[…]
@@ 2 Setup
For our analysis, we use over benchmarks, with their results cutoff at Jan 2024. The benchmarks we used include: AGI Eval (Zhong2023AGIEvalAH), Alpaca (v2) (alpaca_eval), and its length-adjusted version (Dubois2024LengthControlledAA), HuggingFace OpenLLM Leaderboard (open-llm-leaderboard), MMLU (Hendrycks2020MeasuringMM), MAGI (MAGIbenchmark), Chatbot-Arena and MTBench (Zheng2023JudgingLW), Big Bench Hard (suzgun2022challenging). HumanEval (Chen2021EvaluatingLL) ARC (clark2018think), HellaSwag (zellers2019hellaswag), TruthfulQA (lin2022truthfulqa), Winogrande (sakaguchi2019winogrande), GSM8k (cobbe2021traininggsm8k). EQ-Bench (v2) (paech2023eqbench), ArenaHard (arenahard2024) and OpenCompass (2023opencompass). For a wider survey of benchmarks used, see App. 9.1.
Our analysis focuses on evaluating agreement between two benchmarks – a reference benchmark (established and commonly acceptable) and a target benchmark (the one we assess, e.g., a new benchmark). Specifically, agreement is calculated as the correlation over the models ranks (using Kendall (Kendall1938ANM)) or scores (using Pearson (PearsonVIINO)).
We note that an inherent constraint in BAT is the size of the model intersection between the benchmarks (i.e., models appearing in both benchmarks). Benchmarks lacking a sufficiently large set of intersecting models (for this work, we chose ), cannot be reliably used for BAT.
| Recommendations | BAT Variance |
[…]
Reduction | ||||||||
[…]
@@ 3 BAT Methodological Decisions: An Analysis
In the absence of clear guidelines, benchmark creators often make arbitrary choices, without clear justification or consistency across different studies.
@@ 3.1 The Choice of Reference Benchmark Matters
Finding a reference benchmark for BAT is a non-trivial task. One needs to find a well-established benchmark, whose data is readily available, and which exhibits a large enough overlap with the models already evaluated in the target benchmark. Due to the above difficulty, BAT is commonly done against one or two reference benchmarks (Yuan2024PRobELMPR). Benchmarks can be divided into groups according to their measured abilities – for example, holistic benchmarks that aim to measure some loosely-defined construct of overall model quality, such as BigBench (srivastava2023beyond), benchmarks measuring coding abilities (Chen2021EvaluatingLL), math benchmarks (cobbe2021traininggsm8k) etc. Thus, when selecting a reference benchmark, there is often a somewhat arbitrary choice between several possible benchmarks which are all seemingly appropriate.
Figure 2 illustrates the variability stemming from such arbitrary choices: for each target benchmark, different reference benchmarks produce wildly varying agreement scores. For example, Alpaca V2 (second row from above) demonstrates a wide range of agreement levels with other benchmarks, spanning from a mediocre agreement of with MT-bench to a high agreement of with LMSys Arena, even though both of these reference benchmarks are considered to measure similar abilities. This variability calls into question the validity of conclusions based on applying BAT using a single reference benchmark.
To address this issue, we advocate using an aggregated reference benchmark that consolidates multiple benchmarks based on the mean-win-rate; see more on this in §4.
@@ 3.2 The Choice of Models Matters
Figure 5 demonstrates the relationship between the number of models and the variability of BAT results. It shows that with a small amount of models, BAT results can get highly unreliable, with a standard deviation approaching . For instance, in our analysis we found that the Kendall-tau correlation between LMSysArena and MT-Bench can range from approximately 0.65 to 0.99, depending on the particular number of models chosen. Thus, we see that the common practice of using a small number of models for BAT may jeopardise the validity of conclusions.
@@
To quantify this phenomenon, we investigate benchmark agreement where the subset of models selected is not completely random, but is constrained to sets of models that are adjacent in rank (e.g., models 3-7)111Note that the sets of adjacent models were not selected from a specific rank location (e.g., Top, Bottom, Middle) but were randomly selected from the full range. For an analysis of such location-dependent sets, see App 9.2.. Adjacent models have more similar performance. Thus, their score differences and ranking may be less stable, resulting in lower correlation scores. In Figure 3, we show that indeed, for a given number of models, the correlation score when considering adjacent models is lower than that of randomly sampled models, with a stronger effect as the number of models in the subset decreases.
This discrepancy emphasizes the importance of using varying granularities when reporting BAT scores. This would enable managing the expectations of benchmark consumers, who may expect and desire a specific level of granularity (e.g., getting the very best models right, or discriminating between strong and weak models).
3.3 The Choice of Correlation Metric (and Threshold) Matters
@@
Currently, there are no consistent standards for the types and thresholds of correlation metrics. For instance, Liu2021DoQA utilized both rank and score correlations, setting a uniform threshold of for both, whereas sun2023validity exclusively employed rank correlation and opted for a distinct threshold of .
@@ 4 BAT Best Practices
The choice of reference benchmark can hinder the validity of BAT conclusions (§3.1). To circumvent this issue, we propose to combine the results from all benchmarks appropriate for the goal of the BAT (choosing benchmarks that measures similar or dissimilar abilities) into an aggregate reference benchmark, by averaging their model win-rates. For example, when using BAT to validate some efficient holistic benchmark, the reference benchmark should be the aggregate of all available holistic benchmarks. By combining results from a group of benchmarks, the aggregate benchmark can provide both a more stable and robust basis for comparison. Notably, since the aggregate benchmark captures the distribution of relevant results, it constitutes a better measure of the underlying construct represented by the group, called in the literature convergent validity (carlson2012understanding).
@@
BAT based on a small set of models has a large variance (§3.2). We recommend using at least 10 models when performing BAT, preferably more, to ensure more reliable results. These models should represent the entire spectrum of available models, including diverse sizes, architectures, and training methods. Aiming for a random selection ensures equal representation and minimizes bias. Table 1 shows that using this methodology to select models decreases BAT variance by more than %.
@@
Benchmark agreement varies significantly with the range of qualities of the models used in BAT (§3.2), see Figure 1. To manage the expectations of benchmark consumers, who may expect and desire a specific level of granularity (e.g., getting the very best models right, or discriminating between strong and weak models) we recommend reporting agreement scores at multiple resolutions (e.g., 5/10/20 contiguous models, averaging across groups when more models were sampled). This approach provides a more nuanced view on benchmark agreement, highlighting critical distinctions that might otherwise be missed (e.g. the top 3 models are almost never in agreement across benchmarks).
Use a Data-driven Threshold
Using predetermined thresholds to interpret correlation scores may be misleading, as the meaning of "high" or "low" agreement inherently has a relative aspect to it. For instance, consider the correlation scores obtained when using different model granularities. As seen in Figure 3, benchmarks produce very different correlation scores depending on the model subset chosen. Thus two benchmarks that are in low agreement can show high correlation scores in some configurations.
This difficulty calls for the use of a data-driven approach for interpreting agreement scores. Thus, for a given configuration and reference benchmark(s), we compile a distribution of agreement scores of various benchmarks against this reference. Then, we compare the correlation score obtained for the target benchmark with this distribution. Specifically, we calculate the Z-score of the target-reference correlation score relative to this distribution. A target benchmarks that produce a Z-score above are considered in agreement, whereas below it, they are not. As this method incorporates a natural distribution of benchmark agreement scores, it harnesses the natural mean agreement while also incorporating the nature of the distribution, meaning that the notion of agreement becomes relative to the population of existing benchmarks. As more benchmarks are added the test will more accurately reflect the agreement according to the current natural distribution of benchmarks.
@@
[…]
@@
[…]
@@
A user inputs their BAT configuration, including the desired group of reference benchmarks.
@@
[…]
recommends a set of models for evaluation on the target benchmark.
@@
[…]
produces a full BAT report.
[…]

### v2

# Do These LLM Benchmarks Agree? Fixing Benchmark Evaluation with BenchBench

@@ Abstract
Recent advancements in Language Models (LMs) have catalyzed the creation of multiple benchmarks. A crucial task, however, is assessing the validity of the benchmarks themselves. This is most commonly done via Benchmark Agreement Testing (BAT), where new benchmarks are validated against established ones using some agreement metric (e.g., Spearman correlation). Despite the crucial role of BAT for benchmark builders and consumers, there are no standardized procedures for such agreement testing. This deficiency can lead to invalid conclusions, fostering mistrust in benchmarks and upending the ability to choose the appropriate benchmark. By analyzing over 50 prominent benchmarks, we demonstrate how some overlooked methodological choices can significantly influence BAT results, potentially undermining the validity of conclusions. To address these inconsistencies, we propose a set of best practices for BAT and demonstrate how utilizing these methodologies greatly improves BAT robustness and validity. To foster adoption and facilitate future research, we introduce BenchBench111https://github.com/IBM/benchbench, a Python package for BAT, and release the BenchBench-leaderboard222https://hf.co/spaces/ibm/benchbench, a meta-benchmark designed to evaluate benchmarks using their peers.
Do These LLM Benchmarks Agree?
Fixing Benchmark Evaluation with BenchBench
Yotam Perlitz1 Ariel Gera1 Ofir Arviv1 Elron Bandel1 Asaf Yehudai1 Eyal Shnarch1 Michal Shmueli-Scheuer1 Leshem Choshen2,3 1IBM Research AI 2MIT CSAIL 3MIT-IBM {y.perlitz,leshem.choshen}@ibm.com
@@ 1 Introduction
As Language Models (LMs) increasingly excel across a broad range of tasks, new benchmarks – often measuring similar abilities – are constantly proposed. This deluge of benchmarks underscores the importance of Benchmark Agreement Testing (BAT). BAT involves validating a new benchmark by comparing it against established and trusted benchmarks, using statistical agreement metrics. This comparison is based on the performance scores of models across the different benchmarks.
BAT is often used to validate that a new proposed benchmark measures what it was designed to measure. The expectations from this measurement depend on the benchmark’s goal; demonstrating high agreement can serve to show that a new benchmark captures model abilities similar to those measured by established and well trusted benchmarks. (Lei et al., 2023; Viswanathan et al., 2023; Chang et al., 2023; Li et al., 2024b; Prabhu et al., 2024; He et al., 2024). High agreement can also validate that an efficient version of a benchmark (e.g., requiring less compute or labeling) measures the same thing as the original benchmark (Perlitz et al., 2023; Polo et al., 2024; Prabhu et al., 2024; Vivek et al., 2023). In contrast, if a benchmark aims to test a unique trait – one that is not properly covered by existing benchmarks – BAT will be used to demonstrate the disagreement of such benchmarks with existing ones (Yuan et al., 2024; Waldis et al., 2024). The above goals are relevant both for benchmark creators and for benchmark consumers. Creators will typically use BAT to validate the properties of their new benchmark; benchmark consumers might use it to choose which existing benchmark they want to use.
However, despite the wide application of BAT in recent years, there is a glaring absence of common methodology. Specifically, the significance of several methodological decisions in BAT is currently overlooked, undermining the validity of any conclusions made.
@@
We focus on three such critical choices: selecting the reference benchmark (§3.1), the models included in the test (§3.2), as well as the correlation metrics and their interpretation (§3.3). For example, as seen in Figure 2, choosing a different subset of models produces substantially different correlation scores, leading to different conclusions about benchmark agreement. The figure demonstrates that two benchmarks can (and often do) show high agreement across a wide range of models, while agreement over a few top-ranked models remains low.
Building upon our findings, we compile a set of best practices for BAT (§4) and demonstrate their impact (Figure 1 and Table 1). To foster adoption and promote reproducibility, we have implemented these guidelines into BenchBench, a Python package for BAT (§5). BenchBench supplies users not only with a framework but also with the data needed to perform BAT, relieving users of the computational and time burden of gathering multiple benchmarks for comparison. Notably, when using BenchBench, applying our best practices for running BAT will not require further computational resources. Furthermore, BenchBench is built to continually evolve, allowing easy addition of new benchmarks.
Lastly (§5), we introduce the BenchBench-Leaderboard. Using BenchBench as its back-end, the BenchBench-Leaderboard is a dynamic leaderboard that provides easy access to BAT results for established benchmarks. By ranking benchmarks based on their agreement with the user’s desired set of reference benchmarks, the BenchBench-Leaderboard facilitates making informed evaluation decisions.
@@
We release BenchBench, a Python package for BAT implementing the guidelines and incorporating them with the required benchmark data (§5).
@@ 2 Setup
For our analysis, we use over benchmarks, with their results cutoff at Jan 2024. The benchmarks we used include: AGI Eval (Zhong et al., 2023), Alpaca (v2) (Li et al., 2023), and its length-adjusted version (Dubois et al., 2024), HuggingFace OpenLLM Leaderboard (Beeching et al., 2023), MMLU (Hendrycks et al., 2020), MAGI (Paech, 2024), Chatbot-Arena and MTBench (Zheng et al., 2023), Big Bench Hard (Suzgun et al., 2022). HumanEval (Chen et al., 2021) ARC (Clark et al., 2018), HellaSwag (Zellers et al., 2019), TruthfulQA (Lin et al., 2022), Winogrande (Sakaguchi et al., 2019), GSM8k (Cobbe et al., 2021a). EQ-Bench (v2) (Paech, 2023), ArenaHard (Li et al., 2024a) and OpenCompass (Contributors, 2023). For a wider survey of benchmarks used, see App. 9.1.
Our analysis focuses on evaluating agreement between two benchmarks – a reference benchmark (established and commonly acceptable) and a target benchmark (the one we assess, e.g., a new benchmark). Specifically, agreement is calculated as the correlation over the models ranks (using Kendall (Kendall, 1938)) or scores (using Pearson (Pearson, 1895)).
We note that an inherent constraint in BAT is the number of intersecting models between the benchmarks (i.e., models appearing in both benchmarks). Benchmarks lacking a sufficiently large set of intersecting models (for this work, we chose ), cannot be reliably used for BAT.
@@ 3 BAT Methodological Decisions: An Analysis
In the absence of guidelines, benchmark creators often make arbitrary choices, without clear justification or consistency across different studies.
@@ 3.1 The Choice of Reference Benchmark Matters
Finding a reference benchmark for BAT is a non-trivial task. One needs to find a well-established benchmark, whose data is readily available, and which exhibits a large enough overlap with the models already evaluated in the target benchmark. Due to the above difficulty, BAT is commonly done against one or two reference benchmarks (Yuan et al., 2024). Benchmarks can be divided into groups according to their measured abilities – for example, holistic benchmarks that aim to measure some loosely-defined construct of overall model quality, such as BigBench (bench authors, 2023), benchmarks measuring coding abilities (Chen et al., 2021), math benchmarks (Cobbe et al., 2021a), etc. Thus, when selecting a reference benchmark, there is often a somewhat arbitrary choice between several possible benchmarks which are all seemingly appropriate.
Figure 3 illustrates the variability caused by such arbitrary choices: for each target benchmark, different reference benchmarks produce wildly varying agreement scores. For example, Alpaca V2 (second row from above) demonstrates a wide range of agreement levels with other benchmarks, spanning from a mediocre agreement of with MT-bench to a high agreement of with LMSys Arena, even though both of these reference benchmarks are considered to measure similar abilities. This variability calls into question the validity of conclusions based on applying BAT when relying on a single reference benchmark.
To address this issue, we advocate using an aggregated reference benchmark that consolidates results of multiple benchmarks based on the mean-win-rate; see more on this in §4.
@@ 3.2 The Choice of Models Matters
Figure 5 illustrates the relationship between the number of models and the variability of BAT results. It shows that with a small amount of models, BAT results can get highly unreliable, with a standard deviation approaching . For instance, in our analysis we found that the Kendall-tau correlation between LMSysArena and MT-Bench can range from approximately 0.65 to 0.99, depending on the particular number of models chosen. Thus, we see that the common practice of using a small number of models for BAT may jeopardise the validity of conclusions.
@@
To quantify this phenomenon, we investigate benchmark agreement where the subset of models selected is not completely random, but is constrained to sets of models that are adjacent in rank (e.g., models 3-7)333Note that the sets of adjacent models were not selected from a specific rank location (e.g., Top, Bottom, Middle) but were randomly selected from the full range. For an analysis of such location-dependent sets, see App 9.2.. Adjacent models have more similar performance. Thus, their score differences and ranking may be less stable, resulting in lower correlation scores. In Figure 4, we show that indeed, for a given number of models, the correlation score when considering adjacent models is lower than that of randomly sampled models, with a stronger effect as the number of models in the subset decreases.
This discrepancy emphasizes the importance of reporting BAT scores at multiple levels of granularity. This would enable managing the expectations of benchmark consumers, who may expect and desire a specific level of granularity (e.g., getting the very best models right, or discriminating between strong and weak models).
3.3 The Choice of Correlation Metric
(and Threshold) Matters
@@ 3.3 The Choice of Correlation Metric
Currently, there are no consistent standards for the types and thresholds of correlation metrics. For instance, Liu et al. (2021) utilized both rank and score correlations, setting a uniform threshold of for both, whereas Sun et al. (2023) exclusively employed rank correlation and opted for a distinct threshold of .
@@ 4 BAT Best Practices
The choice of reference benchmark can significantly affect the validity of BAT conclusions, as demonstrated by the variability in agreement scores when different single benchmarks are used as references (§3.1, Figure 3). To mitigate this variability, we propose combining the results from all benchmarks appropriate for the goal of the BAT (e.g., benchmarks measuring similar or dissimilar abilities) into an aggregate reference benchmark by averaging their model win-rates. This approach reduces the influence of outliers and provides a more stable and robust measure of agreement, leading to more reliable conclusions. For example, when using BAT to validate some efficient holistic benchmark, the reference benchmark should be the aggregate of all available holistic benchmarks. By combining results from a group of benchmarks, the aggregate benchmark provides both a more stable and robust basis for comparison. Notably, since the aggregate benchmark captures the distribution of relevant results, it constitutes a better measure of the underlying construct represented by the group, called in the literature convergent validity (Carlson and Herdman, 2012).
@@
Use a Data-driven Threshold
Using predetermined thresholds to interpret correlation scores can be misleading, as the relative nature of “high” or “low” agreement varies depending on the context, such as model granularity (§3.3, Figure 4). A more accurate and context-aware assessment can be achieved by using a data-driven approach that compares the target benchmark’s agreement with a reference benchmark (preferably an aggregate) to the distribution of agreement scores from various other benchmarks against the same reference. The steps of this approach are as follows:
1.
Compile a Distribution: Begin by compiling a distribution of agreement scores from various benchmarks relative to the chosen reference benchmark.
2.
Calculate the Target Benchmark’s Z-Score: Next, compare the target benchmark’s correlation score to this distribution by calculating its Z-score. Indicating how the target benchmark’s agreement compares to that of other benchmarks.
3.
Interpret the Z-Score: Benchmarks with a Z-score above are considered to be in agreement with the reference; those below this threshold are not.
By incorporating the natural distribution of benchmark agreement scores, this method ensures that the assessment of agreement is both context-sensitive and adaptive to changes in the benchmark landscape. Furthermore, as more benchmarks are added, the distribution is updated, making the test increasingly reflective of the current landscape of benchmarks measuring the desired trait.
@@
BAT based on a small set of models tends to have high variance, as shown in Figure 5, where the standard deviation of results can reach with fewer models (§3.2). To reduce this variability and enhance reliability, we recommend using at least 10 models, preferably more. A larger and more diverse sample provides a more representative evaluation, minimizing bias and improving result stability. While increasing the number of models does raise computational costs, our recommendation remains practical, given that most model benchmarks already evaluate a larger number of models. These models should represent the entire spectrum of available models, including diverse sizes, architectures, and training methods. Aiming for a random selection ensures equal representation and minimizes bias. Table 1 shows that using this methodology to select models decreases BAT variance by more than %.
@@
Benchmark agreement varies significantly with the range of model qualities considered, as demonstrated in Figure 2 (§3.2). For instance, agreement can be high across a broad range of models but low among top-ranked models, which can mislead benchmark consumers who seek fine-grained distinctions. To address this, we recommend reporting agreement scores at multiple resolutions (e.g., 5/10/20 contiguous models, averaging across groups when more models were sampled). This practice provides a more nuanced and complete picture, allowing users to make informed decisions based on their specific needs. This approach provides a more nuanced view of benchmark agreement, highlighting critical distinctions that might otherwise be missed (e.g. the top 3 models are almost never in agreement across benchmarks).
@@
Properly performing BAT using the above guidelines is not a trivial task. These methodologies require complex statistical tools, reproducible analysis and mostly, access to a large amount of up-to-date benchmarks data. Recognizing this difficulty, we have implemented our recommended workflow into BenchBench, a Python package for BAT, described below.
@@
5 BenchBench - a Package and Leaderboard
We introduce BenchBench, a package implementing the above guidelines - standardizing the practice of BAT – and holding results of multiple benchmarks for a wide variety of reference benchmark choices. The python package is available in GitHub at: github.com/IBM/benchbench.
@@ 5 BenchBench - a Package and Leaderboard
A user enters their BAT configuration, including the desired group of reference benchmarks.
@@
BenchBench recommends a set of models for evaluation on the target benchmark.
@@
BenchBench produces a full BAT report.
In the default functionality, BenchBench expects a list of model scores over the target benchmark, as well as a desired group of reference benchmarks to compare to. It also offers the functionality of proposing a minimal set of models for evaluation, ensuring fair and unbiased comparisons. While offering flexibility to change the defaults, BenchBench’s BAT report includes several granularities of models. BenchBench standardizes arbitrary decisions that hinder reproducibility, following the best practices proposed here. Lastly, BenchBench offers the user to upload their benchmark results to the BenchBench database, enriching the reference benchmark distribution for future efforts, thereby enhancing BAT reliability without additional computational costs. due to running additional reference benchmarks.
We propose the BenchBench-leaderboard, a new leaderboard designed to rank benchmarks according to their agreement to a desired group of reference benchmarks (see Figure 7). To do so BenchBench ranks all submitted benchmarks by comparable standards.
Since the BenchBench-leaderboard is build on top of the BenchBench package, new benchmarks uploaded to the package will be added to the leaderboard as well. Thus, the benchmark will improve with time, taking into account novel benchmarks and measured model traits.